# Uncertainty in latent representations of variational autoencoders optimized for visual tasks

### v2

@@ Methods
The objective function to learn the parameters $\varphi$ and $\phi$ of the encoder and decoder, respectively, is formulated as the Evidence Lower Bound (ELBO)[34]. This takes the form
$\mathcal{L}_{VAE}(\mathbf{x},\varphi,\phi)=-\mathbb{E}_{q_{\phi}(\mathbf{z}|\mathbf{x})}\left[\log p_{\varphi}(\mathbf{x}|\mathbf{z})\right]+\beta_{1}D_{KL}(q_{\phi}(\mathbf{z}|\mathbf{x})||p(\mathbf{z})),$ (1)
@@
where the first term expresses the reconstruction error, and a second term acts as regularization, penalising differences between the prior distribution and the approximate posterior of the latent variables. The hyperparameter $\beta_{1}$ determines a possible extra prior regularization weight, allowing further computational flexibility during training. The expectation in the first term is not explicitly calculated, but a Monte Carlo approximation is taken with a single sample from the variational posterior. With this approximation, the reconstruction error takes either the form of cross-entropy or mean-squared error between an input and its reconstruction from the single sample for the Bernoulli and normal likelihoods, respectively.
Explaining-away VAE
The EA-VAE incorporates into the standard VAE an additional scalar latent variable $s\in\mathbb{R}$ , which is inferred simultaneously with the other latents, $\mathbf{z}$, in the encoder, but acts multiplicatively in the decoder. That is, the mapping $\mathbf{f}(\mathbf{z})$ in the generative model of a usual VAE is supplemented by $s$ such that the mean of the likelihood in the EA-VAE is obtained as $s\,\mathbf{f}(\mathbf{z})$ (Figure 1e, right).
Consequently, the extension of the latent space yields a recognition model jointly for $\mathbf{z}$ and $s$: $q_{\phi}(\mathbf{z},s|\mathbf{x})$. To easily accommodate different potential probability distribution classes for the variational posterior, this joint variational posterior was broken down into two independent variational posterior components, $q_{\phi}(\mathbf{z}|\mathbf{x})$ and $q_{\psi}(s|\mathbf{x})$. As the scaling variable $s$ is meant to account for joint modulation of intensities across pixels, a positive value is justified. Consequently, a reasonable choice for the prior for $s$ is a Gamma distribution. In such a case the variational posterior was also taken as a Gamma. We note however that we explored several alternatives, as discussed in following sections showing no qualitative difference in the results.
Since $\mathbf{z}$ and $s$ are independent random variables, we write the objective function for the EA-VAE as
$\mathcal{L}_{EA-VAE}(\mathbf{x},\varphi,\phi)=-\mathbb{E}_{q_{\phi}(\mathbf{z}|\mathbf{x})}\left[\log p_{\varphi}(\mathbf{x}|\mathbf{z},s)\right]+\beta_{1}D_{KL}(q_{\phi}(\mathbf{z}|\mathbf{x})||p(\mathbf{z}))+\beta_{2}D_{KL}(q_{\psi}(s|\mathbf{x})||p(s)).$ (2)
Importantly, in the case of mapping $s$ by a softplus or an exponential function, the invertible transformation ensures that the Kullback-Leibler (KL) divergence remains unchanged [35]. For reconstruction, both $\mathbf{z}$ and $s$ are sampled from their respective posteriors before decoding.
VAE for inference on natural images
In this work we focus on inference over gray-scale images. That is, the input $\mathbf{x}$ corresponds here to a vector of intensity values for each of the pixels in an image. We first implemented and trained a Sparse VAE [19] identical to the one described in [20]. To allow for flexible learning, the model is constructed using fully connected layers (Figure S1a). In order to foster sparsity, this model employs a multivariate uncorrelated Laplace distribution for the prior distribution with zero mean and unit scale. To ensure that the KL divergence in the ELBO is analytically tractable, we chose to use the Laplace distribution for the variational posterior too: $q_{\phi}(\mathbf{z}|\mathbf{x})=\textrm{Laplace}(\mathbf{z};\boldsymbol{\mu}^{(\mathbf{x})},\boldsymbol{b}^{(\mathbf{x})})$, with $\boldsymbol{\mu}^{(\mathbf{x})}$ and $\boldsymbol{b}^{(\mathbf{x})}$ being the respective mean and scale parameters. These vectors are computed through amortized inference using the encoder for each specific image $\mathbf{x}$. We apply a standard sum of squared errors reconstruction loss, which corresponds to a normal observation noise assumption. That is, the likelihood for a given $\mathbf{x}$ is parameterized by a multivariate uncorrelated Normal distribution $p_{\varphi}(\mathbf{x}|\mathbf{z})=\mathcal{N}(\mathbf{x};f(\mathbf{x}),\boldsymbol{\sigma}_{obs}^{2})$, with $f(\mathbf{x})$ the mean. Thus, sampling the variational posterior, we obtain a reconstruction $\hat{\mathbf{x}}$ and Equation 1 takes the form of
$\mathcal{L}_{VAE}(\mathbf{x},\varphi,\phi)=$ $\frac{1}{2\sigma_{obs}^{2}}\sum_{i=1}^{M}(\hat{\mathrm{x}}_{i}-\mathrm{x}_{i})^{2}+\frac{M}{2}\log{(2\pi\sigma_{obs}^{2})}$ (3)
$+\beta_{1}\sum_{j=1}^{D}\left(-1+|\mu^{(\mathbf{x})}_{j}|-\log{(b_{j}^{(\mathbf{x})})}+b_{j}^{(\mathbf{x})}\exp{(-\frac{|\mu^{(\mathbf{x})}_{j}|}{b^{(\mathbf{x})}_{j}})}\right).$
It is worth mentioning that in cases when the observation noise is assumed to be drawn from an uncorrelated Normal distribution, then Equation 1 can be rewritten by global multiplication of $2\sigma_{obs}^{2}$ so that the $\beta$ hyperparameters absorbe the assumed observation noise [36], that is
$\beta^{\prime}=2\sigma_{obs}^{2}\beta.$ (4)
Further normalization can as well turn the first term in Equation 1 into the mean squared error (MSE). Selecting a larger $\beta$ as done in $\beta$-VAEs then corresponds to performing inference under an enlarged observation noise, fostering disentanglement by use of an uncorrelated prior.
EA-VAE for inference on natural images
The EA-VAE preserves the previous architecture but incorporates a scalar multiplicative latent variable $s$. For natural images patches, this variable is drawn from any specific selected distribution and represents in our model the contrast of the patch, as in previous work modeling natural image patches employing GSMs[11]. We examined different solutions for the implementation of this new architecture, ranging from a homogeneous latent space to having two separate encoders for $\mathbf{z}$ and $s$ (Figure S1), which variants we expand on below.
Scaling variable modeled as a softplus activated Laplace distribution
The central model presented in the main text aims to include a scaling variable with the smallest possible update to the standard VAE architecture. Here $s$ is one of the regular latents, with the same family of probability distributions for both the prior and the variational posterior, but later transformed by a monotonous, invertible function to the positive half-space, by use of a Softplus function (Figure S1a). We refer to this version of the model as the homogeneous version. The loss function in Equation 2 takes the following form
$\mathcal{L}_{EA-VAE}(\mathbf{x},\varphi,\phi)=$ $\frac{1}{2\sigma_{obs}^{2}}\sum_{i=1}^{M}(\hat{\mathrm{x}}_{i}-\mathrm{x}_{i})^{2}+\frac{M}{2}\log{(2\pi\sigma_{obs}^{2})}$ (5)
$+\beta_{1}\sum_{j=1}^{D-1}\left(-1+|\mu^{(\mathbf{x})}_{j}|-\log{(b_{j}^{(\mathbf{x})})}+b_{j}^{(\mathbf{x})}\exp{(-\frac{|\mu^{(\mathbf{x})}_{j}|}{b^{(\mathbf{x})}_{j}})}\right)$
$+\beta_{2}\left(-1+|\mu^{(\mathbf{x})}_{s}|-\log{(b_{s}^{(\mathbf{x})})}+b_{s}^{(\mathbf{x})}\exp{(-\frac{|\mu^{(\mathbf{x})}_{s}|}{b^{(\mathbf{x})}_{s}})}\right),$
where $\mu_{s}$ and $b_{s}$ are the inferred parameters of the scaling variable posterior.
We note that the resulting distribution (Laplace transformed by Softplus activation) does not have tractable moments. Thus we approximated the true posterior mean with samples in Figure 1h, given that the inferred variance of $s$ was low except below the observation noise, where slightly different samples become similarly close to 0 due to softplus activation (Figure S3). Also note the substitution of $D$ in in the first term of Equation 3 for $D-1$ in Equation 5, since we keep the overall number of latent variables unchanged for a fair comparison.
Beyond the homogeneous implementation of the model, we also explored other implementations, as described below. We note that no qualitative difference was observed in the results, highlighting the robustness of the proposal.
Scaling variable modeled as a Log-normal distribution
@@
Scaling variable modeled as a Gamma distribution
A Gamma distribution can be parametrized in terms of the shape and scale parameters $k$ and $\theta$ respectively. For simplicity, we assumed that posteriors $q_{\psi}(s|\mathbf{x})=\textrm{Gamma}(s;k,\theta^{(\mathbf{x})})$ have a fixed shape parameter $k=2$. Since $k$ is fixed, the posterior’s mean, mode and variance depend only on the inferred scale parameter $\theta$. For the prior distribution, the shape parameter was also fixed at $k=2$ and the scale parameter $\theta=1/\sqrt{2}$ was chosen such that the prior had unit variance $k\theta^{2}=1$. A constant shape parameter $k$ means that the posterior standard deviation is always proportional to the posterior mean. This choice is in line with Weber’s Law of perception [37], where the estimation error of a quantity is a constant fraction of its mean. Architecture details in Figure S1b and Table S1. Under these assumptions, Equation 2 can be expressed similar to Equation 5, with the second KL term based upon Gamma distributions instead of Laplace [38]. The results from this model were analogous to the model presented in the main text (Figure S3).
VAEs and EA-VAEs for classic computer vision domains
In contrast to the natural image models, we explored using a convolutional neural network (CNN) architecture to train models on MNIST, fashionMNIST, and chestMNIST images, since convolutional architectures are more commonly employed in these domains (Figure S1c). These consist of encoders and decoders built of both fully-connected and convolutional layers (Figure S1c). Further, as is standard in the ML field, the variational posterior employed was a multivariate uncorrelated Normal distribution $q_{\phi}(\mathbf{z}|\mathbf{x})=\textrm{Normal}(\mathbf{z};\boldsymbol{\mu}^{(\mathbf{x})},\boldsymbol{\sigma}^{2(\mathbf{x})})$, with $\boldsymbol{\mu}^{(\mathbf{x})}$ and $\boldsymbol{\sigma}^{2(\mathbf{x})}$ the respective mean and variance vectors. In the case of the EA-VAE the variational posterior of the multiplicative variable, $s$, was log-normal: $q_{\psi}(s|\mathbf{x})=\textrm{Log-normal}(s;\nu^{(\mathbf{x})},\xi^{2(\mathbf{x})})$. Priors $p(\mathbf{z})$ and $p(s)$ come from the same family as their respective variational posteriors but with zero mean and diagonal unit variance.
Since the Gaussian pixel noise assumption strongly deviates from actual pixel statistics in the case of MNIST (which consists of black and white pixels), we assumed a bernoulli distribution in the likelihood. Accordingly, the cross-entropy reconstruction loss function was employed. In the case of cMNIST, cFashionMNIST and cChestMNIST contrast augmented datasets, the likelihood was parameterized by a multivariate uncorrelated Normal distribution, therefore resulting in mean squared error reconstruction error. The cost function constructed for the MNIST and ChestMNIST models were
$\mathcal{L}_{VAE}(\mathbf{x},\varphi,\phi)=\sum_{i=1}^{M}-(\mathrm{x}_{i}\log(\hat{\mathrm{x}}_{i})+(1-\mathrm{x}_{i})\log(1-\hat{\mathrm{x}}_{i}))+\beta^{\prime}_{1}\sum_{j=1}^{D}\frac{1}{2}\left(-1+(\mu^{(\mathbf{x})}_{j})^{2}-\log{(\sigma_{j}^{2(\mathbf{x})})}+\sigma_{j}^{2(\mathbf{x})}\right),$ (6)
$\mathcal{L}_{EA-VAE}(\mathbf{x},\varphi,\phi)=$ $\sum_{i=1}^{M}-(\mathrm{x}_{i}\log(\hat{\mathrm{x}}_{i})+(1-\mathrm{x}_{i})\log(1-\hat{\mathrm{x}}_{i}))+\beta^{\prime}_{1}\sum_{j=1}^{D-1}\frac{1}{2}\left(-1+(\mu^{(\mathbf{x})}_{j})^{2}-\log{(\sigma_{j}^{2(\mathbf{x})})}+\sigma_{j}^{2(\mathbf{x})}\right)$ (7)
$+\beta^{\prime}_{2}\frac{1}{2}\left(-1+(\nu^{(\mathbf{x})}_{j})^{2}-\log{(\xi_{j}^{2(\mathbf{x})})}+\xi_{j}^{2(\mathbf{x})}\right),$
Datasets
Natural Images
[…]
MNIST and ChestMNIST datasets
VAEs and EA-VAEs were trained on each domain separately. Each original $28\times 28$ pixelsize image was resized to $32\times 32$ due to the 3 layered convolutional encoder and decoder architecture, and the pixel intensity distribution. The pixel intensities for these canonical datasets are in the range $[0,1]$.
Contrast augmented cMNIST, cChestMNIST and cFashionMNIST datasets
Augmented versions of classic MNIST, ChestMNIST and FashionMNIST datasets were synthetically generated in order to have a rich variety in image contrast. Each image in the original dataset was first subtracted by $0.5$, so the pixel intensities move to the range $[-0.5,0.5]$. Then multiplied by a constant $c$ sampled from a Log-normal distribution, and then intensity was corrupted with a white noise vector $\boldsymbol{\eta}$, sampled from a Gaussian distribution of null mean and $\sigma_{obs}$ standard deviation
@@
This process was repeated such that cMNIST, cChestMNIST and cFashionMNIST datasets are 10 times larger than their original forms. $\sigma_{obs}$ values for each dataset are detailed in Table S2.
Shuffled pixels datasets
@@
Models’ training details
All models were trained in Python with Pytorch, either on a GPU NVIDIA TITAN V with a 64GB RAM memory (PC: 1) or on a NVIDIA GeForce RTX 2080 Ti with a 64GB RAM memory (PC: 2). Each model’s hyperparameters are detailed in Table S1 and Table S2.
@@
For the standard VAE, we achieved the best results with beta-annealing: we set $\beta_{1}=0.01$ for the first 100 epochs, linearly increasing to the final $\beta_{1}=1$ over the next 100 epochs. Regarding Softplus-Laplace and Lognormal-Laplace EA-VAE models on natural images, $\beta_{1}=1$ was fixed during the whole training. To avoid collapse in the initial phase, a stronger regularization was applied on the scaling variable in these models for the course of the first 100 epochs ($\beta_{2}=10$), which linearly decreased to $\beta_{2}=1$ during the next 100 epochs. $\beta_{1}$ hyperparameter for the z-scored Gamma-Laplace model was fixed exploratorily such that the decoder’s reconstructions did not present visibly strong deviations from the original input. Results are robust with respect to changes in these hyperparameters, with no abrupt qualitative changes. For cMNIST, cChestMNIST and cFashionMNIST models, $\beta_{1}$ and $\beta_{2}$ were fixed by Equation 4 given the observation noise in the augmentation process.
@@
All models were trained with early stopping so that the parameters were finally set at the epoch with best performance on the validation images. As a reference, the time taken to achieve each respective best epoch on PC 1 in this setting was of $62.2$ hours for the VAE vs $67$ hours for the EA-VAE trained on Van Hateren with Gamma distribution scaling variable. That is, the improvement in uncertainty representation in the EA-VAE comes at an expense of roughly an additional 10% computational time.
In the case of models trained on MNIST and ChestMNIST, although the EA-VAE has $23\%$ more parameters than the VAE, all models took around one hour to complete the 500 epochs. We note that the use of convolutions in this case drastically reduces training times from the case of natural images employing all fully connected layers.
Evaluation of VAEs and EA-VAEs trained on van Hateren
Latent receptive fields
The receptive field $\mathbf{RF}_{j}$ associated to the $j$-th latent dimension was computed through the Spike Triggered Average (STA) technique on the test dataset for all $j=1,...,1800$. The STA was estimated as
$\mathbf{RF}_{j}=\underset{\left\{\mathbf{x}\right\}}{\textrm{mean}}\,\mu_{j}^{(\mathbf{x})}\mathbf{x},$
where $\mu_{j}^{(\mathbf{x})}$ is the $j$-th component of the mean of $\mathbf{x}$’s inferred posterior $q_{\phi}(\mathbf{z}|\mathbf{x})$.
Latent units were discerned as informative or non-informative according to their contribution to the dataset reconstruction. When computing the relative change in the images reconstructed when all units are active, or when a specific unit is replaced by its mean activation throughout the ensemble, it is seen that the units associated to orientation-sensitive filters are clustered with the most contribution, while the second group barely contribute. All results shown in the main text were computed using only the informative latent dimensions.
Uncertainty representation in the latent space
For this, given an ensemble of patches $\left\{\mathbf{x}\right\}$, we infer the contrast $s^{*(\mathbf{x})}$ for each patch as the across-pixel standard deviation of intensities: $s^{*(\mathbf{x})}=\sigma_{pix}^{(\mathbf{x})}$.
The signal mean is defined as the average distance of the posterior’s mean to the prior’s mean (0 in this case) over all patches that have the same inferred contrast,
$\text{SM}(s)=\underset{\left\{\mathbf{x}|s^{*(\mathbf{x})}=s\right\}}{\textrm{mean}}\left\|\boldsymbol{\mu}^{(\mathbf{x})}\right\|.$
Signal variance is computed as the variance of the posterior’s mean of patches of same contrast, averaged across latent dimensions,
[…]
Finally, noise variance is defined as the posterior’s variance vector $\boldsymbol{\sigma}^{2(\mathbf{x})}$ averaged across latent dimensions and observations of equal inferred contrast $s$,
[…]
$\text{SM}(s_{i})=\underset{\left\{\mathbf{x}|s^{*(\mathbf{x})}\in B_{i}\right\}}{\textrm{mean}}\left\|\boldsymbol{\mu}^{(\mathbf{x})}\right\|$
[…]
The associated errors can be expressed as
$\epsilon_{{}_{\text{SM}}}(s_{i})=\frac{1}{\sqrt{N_{i}}}\sqrt{\underset{\left\{\mathbf{x}|s^{*(\mathbf{x})}\in B_{i}\right\}}{\textrm{Var}}\left\|\boldsymbol{\mu}^{(\mathbf{x})}\right\|}$
[…]
$\epsilon_{{}_{\text{NV}}}(s_{i})=\frac{1}{\sqrt{N_{i}}\sqrt{D}}\sqrt{\underset{\left\{j\cup\mathbf{x}|s^{*(\mathbf{x})}\in B_{i}\right\}}{\textrm{Var}}\boldsymbol{\sigma}^{2(\mathbf{x})}}.$
It is worth mentioning that the natural scenes dataset used in this work has patches with varied values of pixel intensity standard deviation $\sigma_{pix}^{(\mathbf{x})}$. As patches with high contrast live far away from the origin in the in the N-dimensional pixel space (since $s_{pix}^{(\mathbf{x})}=\sqrt{\frac{1}{N-1}\sum_{i=1}^{N}(x_{i}-\bar{x})^{2}}=\sqrt{\frac{1}{N-1}}\left\|\mathbf{x}\right\|$), it is expected that the network will have more difficulty exploring farther regions in the pixel space than the regions where low-contrast patches live. Therefore, in evaluation instances of the Gamma-Laplace model, patches with contrast higher than $\sigma_{pix}=2$ were not considered.
Divisive normalization experiments
To characterize the dependencies between latents of the Variational Autoencoders, we calculated conditional distributions for pairs of latents by analyzing the distribution of posterior means for a large number of natural image patches. Across a large number of randomly sampled filter pairs the variance of the conditional distribution was calculated at four quadrants: two central and two flanking quadrants (Figure 5a), which corresponded to low-level, and high-level activations of the conditioned filter, respectively (Figure 5b).
$z_{i}=\frac{L_{i}}{\sqrt{\sum_{j\neq i}w_{ji}L_{j}^{2}+\sigma^{2}}}.$ (9)
Evaluation of VAEs and EA-VAEs trained on MNIST and ChestMNIST
Uncertainty of latent representations

### v3

@@ 2 Methods
The objective function to learn the parameters of the encoder and decoder is formulated as the Evidence Lower Bound (ELBO)[25]. This takes the form
$\mathcal{L}_{VAE}(\mathbf{x},\phi,\psi)=-\mathbb{E}_{q_{\phi}(\mathbf{z}|\mathbf{x})}\left[\log p_{\psi}(\mathbf{x}|\mathbf{z})\right]+\beta D_{KL}(q_{\phi}(\mathbf{z}|\mathbf{x})||p(\mathbf{z})),$ (1)
where the prior, $p(\mathbf{z})$, is usually a standard normal distribution. The first term expresses the reconstruction error, and the second term acts as regularization, penalizing differences between the prior distribution and the approximate posterior. The hyperparameter $\beta$ determines the prior regularization weight, and is linked to the assumed pixel observation noise $\boldsymbol{\sigma}_{obs}^{2}$ [36].
@@
The expectation in the first term is not explicitly calculated, but a Monte Carlo approximation is taken with a single sample from the variational posterior. With this approximation, the reconstruction error takes either the form of the cross-entropy or quadratic loss between an input and its reconstruction for the Bernoulli and normal likelihoods, respectively.
2.1 The explaining-away variational autoencoder (EA-VAE)
In this paper we propose a novel variational inference architecture, specifically designed to improve uncertainty estimates. To that end, the EA-VAE extends the latent space of standard VAEs with a scalar latent variable $s\in\mathbb{R}^{+}$ that acts multiplicatively in the decoder. That is, the mapping $\mathbf{f}(\mathbf{z})$ in the generative model of a standard VAE is supplemented by $s$ such that the mean of the likelihood in the EA-VAE is obtained as $s\,\mathbf{f}(\mathbf{z})$ (Figure˜1e, right). The scaling variable is inferred simultaneously with the other latents, $\mathbf{z}$. The multiplicative contribution of $s$ to the likelihood implements a division of labor between latents: while traditional latents $\mathbf{z}$ encode a variety of details about the input, $s$ accounts for overall variance in the pixel space. Intuitively, the scaling variable can represent global modulations in an image without changes in the content of the image. This division of labor has the consequence that $s$ is directly in charge of modulating the model’s uncertainty about the input. Large values of $s$ encourage the network to represent the input with high fidelity and certainty, while small values of $s$ encourage the posterior of $\mathbf{z}$ to collapse to the prior, reporting maximal uncertainty (see mathematical derivation in Supplementary Section˜1).
For EA-VAEs, the variational posterior is a joint distribution: $q_{\phi,\chi}(\mathbf{z},s|\mathbf{x})$. A widespread approach for VAEs is to learn a variational posterior that assumes independence among the latents. Accordingly, the joint variational posterior of EA-VAE was broken down into two components, $q_{\phi}(\mathbf{z}|\mathbf{x})$ and $q_{\chi}(s|\mathbf{x})$, where $\phi$ and $\chi$ denote the parameters of the two components of the encoder. As $s$ is meant to account for the joint modulation of pixels intensities, a positive value is justified. As shown later, the results are robust against the specific choice of the variational posterior of $s$. We note that whenever a VAE was contrasted with an EA-VAE, the total number of latent variables was preserved to ensure a fair comparison.
The extended generative model yields an updated training objective, that is, an updated ELBO:
$\mathcal{L}_{EA-VAE}(\mathbf{x},\phi,\psi,\chi)=-\mathbb{E}_{q_{\phi}(\mathbf{z}|\mathbf{x})}\left[\log p_{\psi}(\mathbf{x}|\mathbf{z},s)\right]+\beta_{1}D_{KL}(q_{\phi}(\mathbf{z}|\mathbf{x})||p(\mathbf{z}))+\beta_{2}D_{KL}(q_{\chi}(s|\mathbf{x})||p(s)).$ (2)
Here $\beta_{1}$ and $\beta_{2}$ denote the relative strengths of the prior regularization terms of $\mathbf{z}$ and $s$ and respectively. For reconstruction, both $\mathbf{z}$ and $s$ are sampled from their respective posteriors before decoding.
Next, we present the specific VAE and EA-VAE architectures investigated in the paper.
2.2 VAEs for natural images
We explored the contribution of the scaling variable to inference in VAEs using an architecture that featured a linear generative model. In this model the scaling variable, $s$, can be studied in detail as it has a well-defined role for the representation of contrast. Our analysis is aligned with earlier accounts by adopting a large-dimensional latent space, fully connected layers, and a Laplace prior [13, 6]. We trained the model on gray-scale natural image patches.
The Laplace prior serves a sparse coding approach, which has historically been taken to capture natural image statistics and recover the receptive field properties of neurons in the primary visual cortex [30]. We hence employ these models as a starting point for our analysis (though we will thoroughly analyze VAEs with Gaussian priors in later sections). To ensure that the KL divergence in the ELBO is analytically tractable, we employed the Laplace distribution for the variational posterior too: $q_{\phi}(\mathbf{z}|\mathbf{x})=\textrm{Laplace}(\mathbf{z};\boldsymbol{\mu}{(\mathbf{x})},\boldsymbol{b}{(\mathbf{x})})$, with $\boldsymbol{\mu}{(\mathbf{x})}$ and $\boldsymbol{b}{(\mathbf{x})}$ being the respective mean and scale parameters. The generative model for a given $\mathbf{x}$ is parameterized by a multivariate uncorrelated Normal distribution $p_{\psi}(\mathbf{x}|\mathbf{z})=\mathcal{N}(\mathbf{x};\mathbf{f}(\mathbf{z}),\boldsymbol{\sigma}_{\text{obs}}^{2})$, such that $\mathbf{f}(\mathbf{z})=\mathbf{A}\cdot\mathbf{z}$ is a linear transformation, and $\sigma_{\text{obs}}$ is the observation noise (see Supplementary Section˜2.1 for detailed cost function)
The EA-VAE extension preserves the previous model structure and incorporates the scalar multiplicative latent variable $s$ (cf. Figure˜S1 a and b). To avoid degeneracy in the image reconstruction we assume $s$ to be positive. We implemented a range of variants for this: a softplus activated Laplace distribution, a Log-normal distribution, and a Gamma distribution (see Supplementary Section˜2.1 for detailed cost functions). We used the first variant as a default and used the variants as controls.
2.3 VAEs for other classic computer vision domains
We studied the properties of EA-VAEs on standard computer vision datasets (MNIST, ChestMNIST, FashionMNIST, NIH-ChestXray14, Imagenet, see details in Section˜3.1). The corresponding variational autoencoders were not restricted to linear decoders and, as is common practice, convolutional layers were also considered. As is standard in the field, the variational posterior was a multivariate uncorrelated Normal distribution $q_{\phi}(\mathbf{z}|\mathbf{x})=\textrm{Normal}(\mathbf{z};\boldsymbol{\mu}{(\mathbf{x})},\boldsymbol{\sigma}^{2}(\mathbf{x}))$, with $\boldsymbol{\mu}{(\mathbf{x})}$ and $\boldsymbol{\sigma}^{2}{(\mathbf{x})}$ the respective mean and variance vectors. In the case of the EA-VAE the variational posterior of the multiplicative variable, $s$, was log-normal: $q_{\chi}(s|\mathbf{x})=\textrm{Log-normal}(s;\nu{(\mathbf{x})},\xi^{2}{(\mathbf{x})})$. Priors $p(\mathbf{z})$ and $p(s)$ come from the same family as their respective variational posteriors but with zero mean and diagonal unit variance.
When studying the models on the MNIST dataset, we assumed a Bernoulli distribution in the decoder. This choice is justified by the largely black-and-white distribution of pixel intensities of digits. For other datasets (see details in Section˜3.1), the decoder was parameterized by a multivariate uncorrelated Normal distribution (Supplementary Section˜2.1 for detailed cost functions).
3 Experimental Setup
3.1 Datasets
3.1.1 Natural Images
@@
3.1.2 NIST datasets
We employed three popular NIST datasets: MNIST[26], ChestMNIST[46] and FashionMNIST[47], which contain standardized grayscale images for digits, chest radiographs and fashion items, respectively. Each original $28\times 28$ pixelsize image was resized to $32\times 32$ due to the specific used architecture (Section˜3.2). The pixel intensities for these canonical datasets are in the range $[0,1]$.
3.1.3 Contrast augmented NIST datasets
Augmented versions of classic MNIST, ChestMNIST and FashionMNIST datasets were synthetically generated in order to have a rich variety in image contrast. We refer to these datasets as cMNIST, cChestMNIST and cFashionMNIST, respectively. Each image in the original dataset was first subtracted by $0.5$, so the pixel intensities move to the range [0.5,0.5][-0.5,0.5]. Then multiplied by a constant $c$ sampled from a Log-normal distribution, and then intensity was corrupted with a white noise vector $\boldsymbol{\eta}$, sampled from a Gaussian distribution of null mean and $\sigma_{obs}$ standard deviation
@@ 3.1.3 Contrast augmented NIST datasets
This process was repeated such that cMNIST, cChestMNIST and cFashionMNIST datasets resulting in a 10-fold increase in the size of the dataset. $\sigma_{obs}$ values for each dataset are detailed in Table S2.
3.1.4 Shuffled pixels datasets
@@
3.1.5 Higher-resolution datasets
To evaluate uncertainty representations in more complex settings we resorted to two further datasets. Firstly, Imagenet[8, 35], a widely used benchmark for image classification, from which we selected all three categories of human images from the standard validation set: ‘Ball Player’, ‘Groom’ and ‘Scuba Diver’, as well as three object categories: ‘Desk’, ‘Wall Clock’ and ‘Bookcase’. Secondly, we employed the NIH-ChestXray14 dataset[46]. This is a large-scale public dataset of frontal chest radiographs released by the U.S. National Institutes of Health. It is composed of 112,120 images from 30,805 unique patients. The official dataset was split into 70% to train, 10% for validation and 20% to test. Images were resized from $1024\times 1024$ into $224\times 224$, which is standard processing in medical image analysis. Same as NIST datasets, the pixel intensities for these images were set to the range $[0,1]$.
3.2 Implementation and training details
All results presented in the main text correspond to models that were trained in Python with Pytorch 2.x. All models were trained through Pytorch’s Adam optimizer, with learning rate (lr) and weight decay (wd) as detailed in Supplementary Table S1 and Table S2, and all other Adam parameters set to Pytorch’s default. We applied early stopping so that the parameters were finally set at the epoch with best performance on the validation images.
@@ 3.2 Implementation and training details
Following[7], beta-annealing was considered to train models on natural image patches: we set $\beta_{1}=0.01$ for the first 100 epochs, linearly increasing to the final $\beta_{1}=1$ over the next 100 epochs. We observed that in the case of the EA-VAE beta annealing had little effect in the training outcome. The final value $\beta_{1}=1$ was determined exploratorily such that the decoder’s reconstructions did not present visibly strong deviations from the original input. Results are robust with respect to changes in these hyperparameters, with no abrupt qualitative changes.
To rule out the possibility of our findings being contingent on specific architectural choices, we explored three variants of architecture and scaling variable family for the EA-VAE, describing below the central model presented in the main text in which the scaling variable is modeled as a softplus activated Laplace distribution (See control variants details in Supplementary Section˜2.1). The model implements the scaling variable $s$ with minimal update to the standard VAE architecture. Here $s$ is coming from the same family of probability distributions of the regular VAE latents, $\mathbf{z}$ but transformed by a monotonous, invertible function to the positive half-space, by use of a Softplus function. The homogeneous formulation of different latent variables motivates a joint encoder for $\mathbf{z}$ and $s$, yielding a joint variational posterior $q_{\phi,\chi}(\mathbf{z},s|\mathbf{x})$. We refer to this version of the model as the homogeneous version (see Figure˜S1a for detailed architecture). To avoid collapse in the initial phase, a stronger regularization was applied on the scaling variable in these models for the course of the first 100 epochs ($\beta_{2}=10$), which linearly decreased to $\beta_{2}=1$ during the next 100 epochs.
To train the VAE and EA-VAE on the NIST and NIH-ChestXray14 datasets, we also incorporated convolutional layers to parametrize the encoder and decoder components, since they are commonly employed in these domains (see Figure˜S1c and Figure˜S1d for detailed architectures). In this case we used latent dimension $D$ much lower to input size, as a classic dimensionality reduction model (See Supplementary Table S2 for complete hyperparameter values).
For the contrast augmented cMNIST, cChestMNIST and cFashionMNIST models, the architectures were the same as NIST datasets, except for the last layer which lacks the sigmoid activation, corresponding to mean square reconstruction error (see Figure˜S1c for detailed architecture). $\beta_{1}$ and $\beta_{2}$ were fixed by Equation˜S26 given the observation noise in the augmentation process (See Supplementary Table S2).
@@
All models were trained either on a GPU NVIDIA TITAN V with a 64GB RAM memory (PC: 1) or on a NVIDIA GeForce RTX 2080 Ti with a 64GB RAM memory (PC: 2). As a reference, the time taken to achieve each respective best epoch on PC 1 in this setting was of $62.2$ hours for the VAE vs $67$ hours for the EA-VAE trained on Van Hateren natural image patches. That is, the improvement in uncertainty representation in the EA-VAE comes at an expense of roughly an additional 10% computational time. In the case of the models trained on MNIST and ChestMNIST employing convolutional layers, all models took around one hour to complete the 500 epochs.
3.3 Post-training evaluation of VAEs and EA-VAEs